# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 2.4 Synthesis of Prior Work and Remaining Gaps
In reviewing the landscape, it is clear that prior work has made significant strides. LLMs have proven highly effective at classification tasks, such as achieving a balanced accuracy of 94.8% for depression detection and 96.2% for PTSD when incorporated via decision-level fusion in multimodal systems [25]. They have also demonstrated utility in administrative roles and have shown surprising capabilities in generating empathetic responses. These successes, however, also illuminate what is missing. The field’s reliance on datasets derived from social media, while useful for scale, lacks the clinical rigor required for high-stakes applications. Evaluation has often been limited to simple classification or sentiment analysis, failing to assess the complex, multi-step reasoning that defines clinical practice. Even with more advanced methods like LMIQ, challenges of generalization and the risk of hallucination persist.
@@
This analysis reveals a critical gap in the literature: a comprehensive benchmark that is simultaneously grounded in authoritative, expert-validated knowledge and designed to evaluate the full spectrum of complex clinical reasoning. There is an urgent need to move beyond binary classification and simple QA to assess an LLM’s ability to perform tasks that require nuanced understanding, such as formulating a differential diagnosis, creating a multi-stage management plan, or engaging in sequential, case-based reasoning.
@@
The present study is positioned to address these limitations directly. By constructing a novel dataset meticulously curated from the foundational psychiatric textbooks described above, we ensure clinical validity and accuracy. By designing eleven distinct and complex QA task types ranging from diagnostic reasoning to treatment planning and sequential analysis, we create a framework to evaluate the sophisticated cognitive functions required in psychiatry. This work, therefore, moves beyond testing what LLMs can do on simplified tasks and begins to assess their readiness for the complex, rigorous, high-stakes reality of clinical mental healthcare.
2.5 Benchmarks for Evaluating LLMs in Psychiatry and Mental Health
@@
Recently, two notable benchmarks illustrate different approaches for mental health and psychiatry. From Empathy to Action: Benchmarking LLMs in Mental Health with MentalBench-10 and a Novel Cognitive-Affective Evaluation Approach [26] proposed a framework that emphasizes empathy, affective alignment, and the capacity of LLMs to translate understanding into actionable guidance. MentalBench-10 focuses on cognitive-affective evaluations, combining both synthetic and user-generated dialogues to assess empathy, appropriateness of advice, and general mental health support. While this contributes valuable insights into the emotional and supportive dimensions of LLM performance, it primarily emphasizes conversational empathy and cognitive-affective alignment rather than the depth of psychiatric reasoning required for clinical tasks.
Similarly, PsychBench: A comprehensive and professional benchmark for evaluating the performance of LLM-assisted psychiatric clinical practice [27] introduced a wide-ranging evaluation platform that incorporates tasks from diagnosis and treatment to professional communication. Its evaluation metrics include ICD Primary Diagnosis Accuracy, Top Choice Alignment Score (TCAS), Medication Match Score (MMS), and Recommendation Coverage Rate (RCR). Although PsychBench benefits from real-world, diverse clinical cases and validation by psychiatrists, its relatively small dataset (300 cases) restricts generalization, and it still relies partly on synthetic dialogue construction, with limited grounding in authoritative, expert-validated psychiatric sources.
@@
Finally, PsychiatryBench is curated from authoritative psychiatric textbooks and casebooks, aiming to provide comprehensive psychiatric reasoning and knowledge evaluation. It introduces eleven distinct QA tasks including diagnostic reasoning, treatment planning, sequential case analysis, and longitudinal follow-up covering more than 5,300 expert-annotated items. The benchmark relies on LLM-as-judge similarity scoring but is built on an expert-validated, broad knowledge base. By grounding its dataset exclusively in clinically validated sources (e.g., DSM-5-TR Clinical Cases, Stahl’s Essential Psychopharmacology), PsychiatryBench directly addresses the gaps left by prior efforts. In contrast to MentalBench-10, which emphasizes affective alignment, and PsychBench, which provides broad but synthetic coverage, our work offers a benchmark that is simultaneously clinically rigorous, multi-task, and extensible.
3 Datasets
@@
The data set used in this study was manually curated to address the unique challenges of psychiatric question answering (QA), with a focus on a diverse range of clinically relevant tasks. Sourced from authoritative psychiatry textbooks and expert-validated clinical resources, the dataset comprises natural language questions paired with expert-formulated answers. These questions are grounded in real-world psychiatric scenarios. They are designed to evaluate the reasoning, decision-making, and knowledge application abilities of LLMs in both clinical and educational contexts.
@@
Figure 1 provides a visual summary of the composition across eleven task categories, while Table 1 outlines the quantitative distribution of samples per task type. Together, they offer a comprehensive overview of the dataset structure and task diversity used in model evaluation.
@@
The manual collection process ensured that the dataset maintains high-quality, medically accurate content, aligning with standardized diagnostic frameworks such as the DSM-5 and ICD-10, as well as clinical practice guidelines from organizations like NICE and the American Psychiatric Association (APA).
The dataset is structured around eleven distinct QA task types, each representing a different cognitive or clinical function critical to psychiatric practice:
  1. 1.
    Diagnosis: This subset contains 467 expert-annotated question-answer pairs centered on diagnostic reasoning. Each item includes a clinical vignette and a diagnostic query, requiring inference of the most appropriate psychiatric diagnosis. The dataset represents 123 unique diagnostic questions and 295 distinct clinical histories. Examples are drawn from six authoritative sources, including 100 Cases in Psychiatry, Geriatric Psychiatry, and Case Files Psychiatry.
  2. 2.
    Treatment: This subset comprises 258 expert-annotated clinical question-answer pairs focused on treatment decision-making in psychiatry. Each example includes a detailed case history, a mental state examination, and a treatment-related clinical question, accompanied by a model answer curated from authoritative psychiatry textbooks. The scenarios reflect real-world psychiatric care and cover a wide range of disorders and treatment modalities, including pharmacological management, psychological therapies (e.g., CBT, psychodynamic therapy), and patient-specific treatment considerations. Most cases are sourced from well-established educational resources such as 100 Cases in Psychiatry, Case Files Psychiatry, and Geriatric Psychiatry.
  3. 3.
    Treatment Follow-Up: This dataset comprises 27 clinical question-answer pairs focused on the follow-up phase of psychiatric treatment. Each scenario includes a psychiatric case history and a primary follow-up question, with a subset (6 entries) extending to a second-tier follow-up. These items evaluate a model’s capacity to support longitudinal clinical care such as medication adjustments, relapse monitoring, and continued patient engagement. All entries were curated from expert-reviewed psychiatric cases to reflect authentic post-treatment scenarios.
  4. 4.
    Classification: This component includes 418 psychiatric case descriptions, each annotated with both a broad diagnostic category and a specific disorder label. The classification task requires models to interpret narrative clinical data and assign appropriate labels by frameworks such as the DSM-5. Cases span multiple psychiatric disorders and are sourced from various clinical and educational texts. This task is crucial for evaluating a model’s ability to operationalize diagnostic taxonomy and differentiate between related disorders.
  5. 5.
    Management Plan: This dataset consists of 337 case-based question-answer pairs that test a model’s ability to generate psychiatric management plans. Each scenario includes a clinical context, often supplemented with mental and physical examination findings. The goal is to evaluate how models handle care coordination, safety planning, treatment adherence, and multidisciplinary strategies. Questions are sourced from authoritative psychiatric references and reflect decisions clinicians make beyond diagnosis or single-treatment choices.
  6. 6.
    Clinical Approach: This subset contains 56 expertly crafted scenarios aimed at evaluating a model’s ability to reason through complex psychiatric cases. The questions focus on the rationale behind clinical decisions such as ruling out differential diagnoses, identifying red flags, or determining the most appropriate investigative or management step. Each case includes multiple LLM-generated responses, enabling comparative analysis. The dataset emphasizes the "how" of clinical reasoning, offering a rigorous benchmark for evaluating the interpretive depth of psychiatric decision support systems.
  7. 7.
    Mental QA: This dataset includes 326 expert-annotated question-answer pairs that test psychiatric knowledge, definition, and conceptual understanding. Most questions 231 explicitly begin with "define," while others explore variations such as "what is," "differentiate," and "describe." Questions are sourced primarily from the Casefile and Case Vignettes series, and include items such as "How would you define thought withdrawal?" or "What are the key diagnostic clues that suggest this patient has histrionic personality disorder?" This dataset targets factual psychiatric knowledge and terminology aligned with DSM-5, ICD-10, and formal psychiatry training curricula.
  8. 8.
    Sequential Question Answering: This dataset comprises 32 clinical vignettes, each followed by a structured sequence of five diagnostic and management questions. The questions reflect a progressive clinical reasoning path, covering: (1) differential diagnosis, (2) supporting evidence and additional needed information, (3) etiological factors, (4) treatment options, and (5) prognosis. The dataset supports multi-turn reasoning and longitudinal clinical understanding. Each case is also tagged with a classification label (32 unique categories), including disorders like schizophrenia, psychotic depression, and substance-related disorders. This structure is ideal for evaluating retention, reasoning, and contextual depth in LLMs.
  9. 9.
    Multiple Choice Questions (MCQs):  This task group encompasses structured assessment formats commonly used in psychiatric exams. It is subdivided into four types of multiple-choice datasets:
    • MCQ:
      1473 standard multiple-choice questions, each with 4-5 answer options. These are clinical knowledge checks grounded in patient histories and symptom-based reasoning. For instance, one item describes a 40-year-old male with choreiform movements and a family history of psychiatric illness, requiring identification of the gene location associated with Huntington’s disease.
    • Exams:
      A curated exam-style dataset, designed to simulate standardized testing disorders. This set is split into two parts for evaluation: EMI 300 and MCQ 300. These items assess factual recall and decision-making across psychiatric domains.
    • EMI (Extended Matching Items):
      A clinical reasoning task where a thematic list of 8-15 options is presented, followed by multiple clinical vignettes requiring matched answers. The EMI format encourages fine-grained differentiation across closely related disorders or treatments. The dataset contains 277 whole EMI clusters, each with shared headers and grouped questions.
    • EMI Separated:
      To support fine-tuning and task-specific evaluation, the full EMI format was disaggregated into 1037 standalone question items. Each retains its original header and option list for context, but functions as an independent unit. This format simplifies scoring and improves input granularity during training.
      Although the original data includes answer explanations, these were excluded in modeling tasks to maintain fairness and prevent leakage.
Each task type is intended to test different facets of psychiatric competence, ranging from factual recall and concept definitions to complex clinical reasoning and longitudinal patient care. This diversity in task design allows for a comprehensive evaluation of both general-purpose and domain-specialized LLMs.
Table 1 presents a matrix of the dataset. Each row corresponds to a specific educational resource (book), and each column represents one of the eleven QA tasks. A checkmark (✓) indicates that the book contains content suitable for that task. For example, Case Files Psychiatry supports nearly all tasks, making it ideal for multi-task learning models. Conversely, MCQs in Psychology focus solely on multiple-choice questions, and the Casebook in Abnormal Psychology is annotated for its relevance to follow-up care and classification.
This matrix provides a clear taxonomy of task-resource alignment, allowing for filtered dataset construction. Researchers can, for example, extract subsets specific to diagnostic reasoning or build MCQ-only datasets for exam simulation. The manual annotation process guarantees transparency, interpretability, and adaptability to various downstream applications.
The curated nature of this dataset, grounded in expert-reviewed content, ensures its relevance to real-world psychiatric practice. It offers a robust and modular benchmark for evaluating LLMs in clinical decision support, medical training, and knowledge-grounded natural language processing. By incorporating diverse task types and detailed evaluation protocols, this dataset provides a comprehensive framework for measuring model performance and guiding improvements in mental health AI.
@@
Table 1: The table shows which psychiatry textbooks and case-based resources include content related to specific clinical tasks. A check mark (✅) indicates coverage of a task, helping identify which resources contribute to each area in the dataset collection.
@@
Books Diagnoses Treatment Treatment Follow-Up Classification Management Plan Clinical Approach Mental QA Sequential QA MCQs
Geriatric psychiatry [28]
Case Files Psychiatry [29]
Clinical Psychology Casebook [30]
DSM-5 Clinical Cases [31]
DSM-5-TR® Self-Exam Questions[32]
100 Cases in Psychiatry [33]
Core Clinical Cases [34]
Casebook in Abnormal Psychology [35]
Revision MCQ and EMI [36]
Psychiatry case vignettes [37]
Clinical Cases in Psychiatry [38]
Stahl’s Psychopharmacology Cases [39]
Stahl’s Essential Cases [40]
MCQs in Psychology [41]
@@
Multiple Choice Questions 1473
@@
Total 5308
4 Experimental Setup
This section outlines the experimental framework used to evaluate the performance of LLMs across a diverse set of psychiatric question-answering (QA) tasks. We describe the eleven task types designed to test specific aspects of psychiatric reasoning and detail the models employed, including both general-purpose LLMs and domain-specialized medical models. The evaluation methodology includes both automatic metrics and LLM-as-a-judge assessments to capture performance on accuracy, reasoning depth, and clinical alignment. By structuring our experiments around realistic clinical tasks and rigorous evaluation criteria, we aim to produce insights into the capabilities and limitations of current LLMs in mental health domains.
4.1 Task Types
The task addressed in this work is a Question Answering (QA) problem in the medical domain, with a specific focus on psychiatry. The dataset consists of natural language questions and expert-curated answers derived from authoritative psychiatry textbooks. These questions span a variety of clinically relevant scenarios and concepts, reflecting the types of reasoning and decision-making tasks encountered in psychiatric practice and education.
Unlike traditional QA systems focused on a single type of question, this dataset encompasses eleven distinct QA tasks, each representing a different cognitive or clinical function in psychiatry. These tasks range from diagnostic reasoning to treatment planning, classification, and sequential case-based analysis. The goal is to evaluate the ability of QA models to understand, reason, and generate medically accurate responses across these varied task types.
Given the complexity and depth of psychiatric knowledge, this work aligns with applications in clinical decision support, medical education, and knowledge-grounded language modeling. The QA challenges posed by this dataset require both factual recall and higher-order reasoning consistent with medical standards such as the DSM-5 and ICD-10.
Below, we briefly describe each of the eleven task types included in the dataset:
4.1.1 Task 1: Diagnoses

### v2

@@ 2.4 Synthesis of Prior Work and Remaining Gaps
In reviewing the current landscape, prior work has made significant strides in applying LLMs to psychiatric and mental health contexts. LLMs have proven highly effective at classification tasks, such as achieving a balanced accuracy of 94.8% for depression detection and 96.2% for PTSD when incorporated via decision-level fusion in multimodal systems[hassan2025leveragingembeddingtechniquesmultimodal]. They have also demonstrated utility in administrative and workflow-support roles within healthcare settings[Tripathi2024_JAMIA_AdminTasks], and have shown surprising capabilities in generating empathetic responses comparable to, and sometimes exceeding, those of human clinicians[Ayers2023_JAMA_ChatbotEmpathy, Sorin2024_JMIR_EmpathyReview].
These advances, however, illuminate persistent methodological and conceptual gaps. Much of the field’s progress relies on datasets derived from social media platforms, which, while useful for large-scale data acquisition, lack the clinical rigor required for high-stakes psychiatric applications[Rai2024_PNAS_RaceMarkers, Guo2024_MentalHealthLLMsReview]. Evaluation approaches have often been restricted to simple classification or sentiment analysis tasks, which fail to capture the multi-step reasoning and clinical judgment that define psychiatric practice[Tam2024_npjDM_QUEST, Bedi2024_JAMA_SystematicReview, Chen2025_EvalLLMsAgentsHealthcare]. Even with more advanced reasoning frameworks, such as those used in multimodal IQ-style benchmarks, persistent challenges remain regarding generalization and the risk of hallucination in model outputs[Bang2025_ACL_HalluLens, Kalai2025_OpenAI_WhyHallucinate, deHond2024_LancetDH_Validation].
@@
This synthesis reveals a critical gap in the literature: the absence of a comprehensive benchmark that is both grounded in authoritative, expert-validated psychiatric knowledge and designed to evaluate the full spectrum of complex clinical reasoning. There is an urgent need to move beyond binary classification and simple question-answering tasks to assess an LLM’s ability to perform sophisticated, clinically relevant reasoning such as formulating differential diagnoses, developing multi-stage treatment plans, and engaging in sequential, case-based analysis.
@@
The present study is designed to address these limitations directly. By constructing a novel dataset meticulously curated from foundational psychiatric textbooks[Barnhill2022_DSM5TRClinicalCases] we ensure clinical validity, safety, and alignment with professional standards. Furthermore, by introducing eleven distinct question-answer (QA) task types ranging from diagnostic reasoning to treatment planning and sequential analysis, this framework provides a rigorous, multi-dimensional evaluation of LLM performance in psychiatric contexts. In doing so, it moves the field beyond surface-level competence toward assessing true readiness for the complex, high-stakes reality of clinical mental healthcare.
2.5 Benchmarks for Evaluating LLMs in Mental Health
@@
Recently, two notable benchmarks illustrate different approaches for mental health and psychiatry. From Empathy to Action: Benchmarking LLMs in Mental Health with MentalBench-10 and a Novel Cognitive-Affective Evaluation Approach [anonymous2025from] proposed a framework that emphasizes empathy, affective alignment, and the capacity of LLMs to translate understanding into actionable guidance. MentalBench-10 focuses on cognitive-affective evaluations, combining both synthetic and user-generated dialogues to assess empathy, appropriateness of advice, and general mental health support. While this contributes valuable insights into the emotional and supportive dimensions of LLM performance, it primarily emphasizes conversational empathy and cognitive-affective alignment rather than the depth of diagnostic reasoning. Similarly, PsychBench: A comprehensive and professional benchmark for evaluating the performance of LLM-assisted psychiatric clinical practice [liu2025psychbenchcomprehensiveprofessionalbenchmark] introduced a wide-ranging evaluation platform that incorporates tasks from diagnosis and treatment to professional communication. Its evaluation metrics include ICD Primary Diagnosis Accuracy, Top Choice Alignment Score (TCAS), Medication Match Score (MMS), and Recommendation Coverage Rate (RCR). Although PsychBench benefits from real-world, diverse clinical cases and validation by psychiatrists, its relatively small dataset (300 cases) restricts generalization, and it still relies partly on synthetic dialogue construction, with limited grounding in authoritative, expert-validated psychiatric sources.
@@
In contrast to these existing efforts, our proposed benchmark, PsychiatryBench, represents a novel contribution that emphasizes clinically grounded, expert-validated evaluation. PsychiatryBench is curated from authoritative psychiatric textbooks and casebooks, aiming to provide comprehensive diagnostic reasoning and knowledge evaluation. It introduces eleven distinct QA tasks including diagnostic reasoning, treatment planning, sequential case analysis, and longitudinal follow-up covering more than 5,188 expert-annotated items. The benchmark relies on LLM-as-judge similarity scoring but is built on an expert-validated, broad knowledge base. By grounding its dataset exclusively in clinically validated sources (e.g., DSM-5-TR Clinical Cases, Stahl’s Essential Psychopharmacology), PsychiatryBench directly addresses the gaps left by prior efforts. In contrast to MentalBench-10, which emphasizes affective alignment, and PsychBench, which provides broad but synthetic coverage, PsychiatryBench advances the field by offering a benchmark that is simultaneously clinically rigorous, multi-task, and extensible.
3 The PsychiatryBench Benchmark: Dataset Curation and Task Design
@@
The dataset created in this study was manually curated to address the unique challenges of psychiatric reasoning, with a focus on a diverse range of clinically relevant tasks. Sourced from authoritative psychiatry textbooks and expert-validated clinical resources, the dataset comprises natural language questions paired with expert-formulated answers. These questions are grounded in real-world psychiatric scenarios. They are designed to evaluate the reasoning, decision-making, and knowledge application abilities of LLMs in both clinical and educational contexts. It is important to clarify the benchmark’s intended clinical setting and acuity level. PsychiatryBench is primarily designed as a research and pedagogical benchmark to support model development and the systematic evaluation of psychiatric reasoning tasks. As such, its current focus is on adult outpatient psychiatric reasoning and not severely acute clinical contexts.
This emphasis on outpatient scenarios, where structured diagnostic and treatment reasoning can be most effectively assessed, means that the benchmark does not comprehensively represent high-acuity, emergency, or inpatient psychiatry scenarios (such as acute mania requiring admission, delirium, or imminent suicidality). We acknowledge that these contexts represent distinct and critical aspects of clinical care. This focus is a deliberate scoping choice, and the exclusion of these high-acuity presentations is further discussed as a limitation in Section 6.
@@
Figure 1 provides a visual summary of the composition across task categories, while Table 1 outlines the quantitative distribution of samples per task type. Together, they offer a comprehensive overview of the dataset structure and task diversity used in model evaluation.
@@
The manual collection process ensured that the dataset maintains high-quality, medically accurate content, aligning with standardized diagnostic frameworks such as the DSM-5 and ICD-10, as well as clinical practice guidelines from organizations like the National Institute for Health and Care Excellence (NICE) and the American Psychiatric Association (APA).
3.1 PsychiatryBench Development Pipeline
Our methodology, shown in Figure 2, begins with systematic manual extraction of content from standard evidence-based psychiatry textbooks. Subject matter experts thoroughly review and select clinical material with high educational and diagnostic value, vibrant case studies, descriptive patient scenarios, and disorder-specific treatment information. From this material, a diverse set of question-answer (QA) pairs is constructed to represent real-world diagnostic reasoning. The QA items are crafted across eleven core task types: Diagnosis, Treatment, Mental QA, Sequential QA, Management plan, Treatment Follow-up ,Clinical Approach, Classification, Multiple Choice Questions (MCQ), exams, and Extended Matching Items (EMI).
@@ 3.1 PsychiatryBench Development Pipeline
Table 1: A summary of the psychiatry textbooks and case-based resources used for the dataset, showing the specific tasks covered by each source. A check mark (✅) indicates coverage of a task.
@@
Books Diagnoses Treatment Treatment Follow‐Up Classification Management Plan Clinical Approach Mental QA Sequential QA MCQs
Geriatric psychiatry [hategan2024geriatric]
Case Files Psychiatry [toy2009casefiles-psychiatry]
Clinical Psychology Casebook [choo2019clinical-psychology-casebook]
DSM-5 Clinical Cases [barnhill2013dsm5-clinical-cases]
DSM-5-TR Self-Exam Questions[muskin2023dsm5tr-self-exam]
100 Cases in Psychiatry [wright2017_100casespsychiatry]
Core Clinical Cases [clark2011_core-clinical-cases-psychiatry]
Casebook in Abnormal Psychology [brown2016_casebook-abnormal-psychology]
Revision MCQ and EMI [puri2011_revision-MCQs-EMIs]
Psychiatry case vignettes [pall2002_underground-vignettes]
Clinical Cases in Psychiatry [altamura2019_clinical-cases-psychiatry]
Stahl’s Psychopharmacology Cases [cooper2021_stahl-case-studies3]
Stahl’s Essential Cases [radonjic2024_case-studies5]
@@
Multiple Choice Questions 1353
@@
Total 5188
Following construction, the QA items are passed through a rigorous filtering process designed to ensure clarity, clinical validity, and educational utility. Ambiguous, overly simple, or redundant items are removed, and the remaining questions are manually reviewed for content balance and diagnostic diversity. These filtered QA items form the foundation of the PsychiatryBench dataset. In this phase, questions are sorted by task type and formatted consistently, allowing for structured downstream evaluation. A comprehensive data cleaning step is performed next, standardizing answer styles, resolving typographic issues, and ensuring uniform metadata. The result is a refined, high-quality QA dataset that tests not only factual recall but also diagnostic reasoning, treatment planning, and sequential clinical decision-making in psychiatry.
Psychiatry Bench evaluates these dimensions across its eleven task types. We acknowledge that some of our source texts, such as DSM-5-TR Clinical Cases and 100 Cases in Psychiatry, incorporate elements of the biopsychosocial model and functional assessment in their discussions. However, the benchmark tasks primarily emphasize diagnostic accuracy, treatment decision-making, and knowledge application. Notably, we do not systematically assess broader integrative or systemic dimensions such as narrative biopsychosocial formulation (holistic case conceptualization integrating biological, psychological, and social factors), detailed functional assessment, or systems-level collaboration (e.g., multidisciplinary team planning), nor do we assess interactive communication or conversational coherence.
3.2 Clinical Task Design and Implementation:
The overall objective of capturing the multifaceted reasoning processes essential to psychiatric practice served as a guide for the creation and execution of the PsychiatryBench clinical tasks. Each task type was designed to assess a distinct aspect of clinical competency, including sequential reasoning, longitudinal management, treatment decision-making, and diagnostic formulation. Real clinical cases were converted into structured question–answer pairs that emulate real-world diagnostic and therapeutic reasoning, using only reliable, expert-validated psychiatric sources. The inclusion of medical history and mental state examination (MSE) elements varies across cases, reflecting the structure of the original source texts rather than deliberate design choices. This variability was retained to preserve the authenticity and instructional style of the reference materials. This section outlines the conceptual rationale, data generation procedures, and task-specific frameworks used to operationalize these competencies, thereby providing a foundation for the systematic, clinically grounded evaluation of LLMs in psychiatry. For additional transparency, representative sample items from multiple PsychiatryBench task types are provided in Appendix 1.3.
3.2.1 Diagnosis